OpenStudy (anonymous):

(x/x+5)-(5/x+5)=x^2+25/x^2-25

OpenStudy (phi):

to avoid confusion, you should write this as x/(x+5) - 5/(x+5)= (x/x+5) means (technically speaking!) \[\frac{x}{x} +5 \] which is 1+5=6

OpenStudy (anonymous):

sorry

OpenStudy (phi):

btw, are you sure you don't have a typo?

OpenStudy (anonymous):

no

OpenStudy (phi):

the left side is just like your previous problem

OpenStudy (anonymous):

the answer I came out with is there is no solution

OpenStudy (phi):

what happens if you replace everywhere you see x with a zero?